Parth (parthkohli):

I just noticed a common mistake when I was thinking about the answer to the question \(3x^2 = 18x, x = ?\). \(\textbf{Wrong way:}\) Divide both sides by \(3\). \( \color{Black}{\Rightarrow x^2 = 6x }\) Divide both sides by \(x\). \(x = 6\) I admit that it does it for \(6\), but we are missing one solution. \(\textbf{Correct way:}\) Form a quadratic equation from the given: \(\Rightarrow 3x^2 - 18x = 0\) Divide both sides by \(3\). \(\Rightarrow x^2 - 6x = 0\) Factor. \(\Rightarrow x(x - 6) = 0\) Use the zero-product rule now.

Parth (parthkohli):

I hope that this prevented you from committing mistakes in the future.

OpenStudy (anonymous):

You know this now??

Parth (parthkohli):

?

OpenStudy (anonymous):

The degree of x is here 2 that is sufficient to tell that there must be two values of x..

Parth (parthkohli):

Yeah. But that's not exactly what I pointed out here.

Parth (parthkohli):

Also, \(x^2 + 10x + 25 = 0\) has only one solution, so your statement doesn't apply always.

OpenStudy (anonymous):

It has two solutions dear.. The condition is D = 0 here which says that the two solutions are real and equal here.. x = -5 here..

OpenStudy (unklerhaukus):

so i cannot just divide by x when ever i feel like it eh,?

Parth (parthkohli):

In this case - no.

OpenStudy (unklerhaukus):

i can always divide by three though ,

OpenStudy (anonymous):

Yes you can divide by constant but you don't know the value of x because it is a variable so its values can be 0.. So dividing by 0 is always avoided..

Parth (parthkohli):

Yep yep

OpenStudy (unklerhaukus):

ok, i understand now

hero (hero):

Actually Parth, you skipped a step: Factor out the three first: \(3x(x - 6) = 0\) Then, divide by 3: \(x(x-6)=0\) Anytime you divide by a variable, you are eliminating a solution set, so it is definitely an algebraic a "no-no".